# Consistency Flow Model Achieves One-step Denoising Error Correction Codes

### v1

@@ Abstract
Error Correction Codes (ECC) are fundamental to reliable digital communication, yet designing neural decoders that are both accurate and computationally efficient remains challenging. Recent denoising diffusion decoders with transformer backbones achieve state-of-the-art performance, but their iterative sampling limits practicality in low-latency settings. We introduce the Error Correction Consistency Flow Model (ECCFM), an architecture-agnostic training framework for high-fidelity one-step decoding. By casting the reverse denoising process as a Probability Flow Ordinary Differential Equation (PF-ODE) and enforcing smoothness through a differential time regularization, ECCFM learns to map noisy signals along the decoding trajectory directly to the original codeword in a single inference step. Across multiple decoding benchmarks, ECCFM attains lower bit-error rates (BER) than autoregressive and diffusion-based baselines, with notable improvements on longer codes, while delivering inference speeds up from 30x to 100x faster than denoising diffusion decoders.
@@ 1 Introduction
Error Correction Codes (ECC) play a central role in modern digital communications and have been essential in a wide range of applications, including wireless communication and data storage. The core task of an ECC decoder is to recover a message from a received signal corrupted by noise during transmission. Recently, inspired by the great success of deep neural networks in domains such as computer vision [he2016deep], generative modeling [goodfellow2020generative, ho2020denoising, song2020score], and natural language processing [vaswani2017attention, brown2020language], neural network-based ECC decoders were introduced, which have been shown to be capable of improving the performance scores of conventional, problem-specific algorithms such as Belief Propagation (BP)[richardson2002capacity] and Min-Sum (MS)[fossorier1999reduced].
The existing neural decoders can be categorized into two groups. Early model-based approaches [lugosch2017neural, nachmani2019hyper, zhu2020learning, dai2021learning, kwak2022neural] achieved successful results by integrating neural networks into conventional decoding algorithms. However, their reliance on problem-specific structures can limit their applicability as a general-purpose decoder. To address this limitation and train a general-purpose decoder, model-free decoders enable the extension of a general neural bet architecture without prior knowledge of the decoding algorithm. While the early proposals applying fully connected neural net architectures [gruber2017deep, cammerer2017scaling, kim2018communication] could lead to overfitting, the preprocessing techniques that utilize magnitude and syndrome vectors [bennatan2018deep] have been effective in reducing the error. These improvements have led to efficient transformer-based decoders [choukroun2022error], which leverage the self-attention mechanism to improve the numerical performance on short block codes. The auto-regressive method in transformer-based decoders has been improved in several recent works [choukroun2024foundation, choukroun2024learning]. Notably, the recent work by [parkcrossmpt] integrates cross-attention between magnitude and syndrome, resulting in the state-of-the-art performance in the decoding task.
@@
Figure 1: Illustration of our proposed Error Correction Consistency Flow Model (ECCFM). ECCFM learns to map received signals from the trajectories to a single, consistent codeword prediction $\mathbf{x}_{0}$, represented by a $\delta$ function.
@@
In parallel, the remarkable success of diffusion generative models [ho2020denoising, song2020score] across various domains [rombach2022high, he2022latent, lou2023discrete, chen2024videocrafter2, nie2025large] has inspired a new direction for deep learning-based ECC. Diffusion models train a noise estimator to gradually denoise a noisy input data and reverse a forward noise manipulation process, generating high-quality samples by an iterative application of a neural net denoiser to an input Gaussian noise. The DDECC method [choukroun2022denoising] extends the denoising diffusion framework to ECC, naturally modeling the AWGN channel as a forward diffusion process. In this setup, a time-dependent transformer learns to denoise the received signal iteratively, recovering the original codeword. This framework provides complementary gains over auto-regressive methods and establishes a new state-of-the-art in terms of Bit Error Rates (BER), particularly for long codes and low SNR, due to its iterative refinement denoising process.
However, the iterative denoising of DDECC has introduced a new challenge: the multi-step sampling process incurs significant computational overhead and high latency, reducing the practicality of the approach in real-world applications. This motivates the following question: Can a decoder maintain the performance achieved by the denoising diffusion frameworks at a lower latency that meets the latency requirements of several communication settings?
In this work, we aim to address this question by proposing Error Correction Consistency Flow Models (ECCFM), an architecture-agnostic training framework designed to construct high-fidelity, one-step denosing decoders. Inspired by recent advances in consistency models [song2023consistency], we formulate the decoding process as a Probability-Flow Ordinary Differential Equation (PF-ODE), which seeks to learn an optimal trajectory from the noisy signal distribution to the clean codeword distribution as shown in Figure 1. However, a key challenge in the ECC setting is that the noise level cannot be measured by a continuous timestep as in image generation; instead, it is indicated by the sum of syndrome error in [choukroundenoising]. A direct adaptation of consistency models struggles to learn this highly non-smooth decoding trajectory.
To overcome the mentioned challenge, we propose a differential time condition, using a soft-syndrome formulation from [lau2025interplaybeliefpropagationtransformer] to regularize the reverse ODE process. This ensures the decoding trajectory is smooth for single-step mapping. Building upon this theoretical foundation, the ECCFM decoder is trained directly with a consistency objective and a soft-syndrome regularization term, enabling it to map any noisy received signal to the estimated original codeword in a single, efficient inference step.
@@
We demonstrate the effectiveness of ECCFM through several numerical experiments on a diverse set of standard codes, including BCH, Low-Density Parity-Check (LDPC), Polar, MacKay, and CCSDS. The results show that ECCFM can improve Bit-Error-Rate (BER) compared to leading auto-regressive methods, exhibiting particularly strong gains on longer codes with code lengths above 200. Notably, it achieves an inference speedup of 30x to 100x over the existing iterative denoising diffusion methods in a single step, while maintaining comparable decoding performance. In our numerical evaluation, ECCFM offers a model-free training diagram for building ECC decoders that achieves near state-of-the-art performance with low latency, and requires only a single-step inference suited for most real-world applications.
@@ 2 Related Works
Neural Network-based ECC Decoders. Neural network-based decoders are broadly categorized into model-based and model-free approaches. Model-based decoders augment conventional algorithms, such as Belief Propagation (BP)[richardson2002capacity] and Min-Sum (MS)[fossorier1999reduced], by using neural networks to learn the message-passing process. This paradigm has been extensively explored across various architectures and code types [dai2021learning, kwak2022neural, kwak2023boosting, lugosch2017neural, nachmani2019hyper, nachmani2021autoregressive, marinkovic2010performance, zhu2020learning], consistently achieving superior performance over conventional algorithms [matsumine2024recent]. However, they are often limited by challenges in capturing long-range dependencies and the reliance on the underlying decoding algorithm.
@@
In contrast, model-free decoders treat decoding as a learning problem without depending on problem-specific algorithms. While early fully-connected architectures [gruber2017deep, cammerer2017scaling, kim2018communication] struggled with overfitting, [bennatan2018deep] proposes the pre-processing by decomposing the magnitude and syndrome vector to address overfitting issues for model-free decoders. Inspired by the recent breakthrough of Transformers [vaswani2017attention], ECCT [choukroun2022error] pioneered by applying self-attention to the channel output, and modeling decoding as an auto-regressive sequence-to-sequence task. Several works have improved based on this approach: FECCT [choukroun2024foundation] improved generalization, DC-ECCT [choukroun2024learning] enabled joint encoder-decoder training, and CrossMPT [parkcrossmpt] introduced cross-attention to achieve better performance and efficiency among auto-regressive methods. Recently, diffusion generative models have emerged as a powerful alternative. DDECC [choukroundenoising] frames decoding as a denoising process, modeling the AWGN channel as the forward diffusion step, and offers performance gains over auto-regressive decoders.
Diffusion Generative Models. Diffusion generative models [sohl2015deep, ho2020denoising, song2020score, karras2022elucidating] learn to reverse a forward noising process by estimating the data’s score function [stein1972bound]. They have achieved state-of-the-art performance in synthesizing high-fidelity samples across diverse domains like images [dhariwal2021diffusion, rombach2022high, podell2024sdxl], video [he2022latent, blattmann2023align], text generation [lou2023discrete, nie2025large], and graphs [sun2023difusco, li2023t2t, lei2025boosting]. However, a primary limitation of diffusion models is the significant computational overhead during inference, due to the iterative nature of the denoising process. To mitigate this, numerous accelerated sampling methods have been developed [song2020denoising, lu2022dpm].
This framework is particularly well-suited for Error Correction Code (ECC) decoding, where the AWGN channel naturally models the forward process [choukroun2022denoising]. While diffusion-based decoders have reached state-of-the-art performance [choukroun2022denoising, park2024crossmpt], they inherit the same computational inefficiency compared to methods like auto-regressive decoders [choukroun2022error, park2024crossmpt]. To address this issue, we draw inspiration from Consistency Models (CMs) [song2023consistency, song2023improved, geng2024consistency], a novel technique for accelerating diffusion models. CMs are designed to directly learn the reverse denoising trajectory, enabling mapping noisy samples to the target data distribution in one step and maintaining high generation quality.
@@ 3 Preliminaries
The objective of a decoder is to estimate the original codeword $\hat{x}$ from the noisy signal $y$. An essential tool for error detection is the syndrome, calculated from a hard-demodulated formulation of the received signal, $y_{b}$, performed as $y_{b}=\text{bin}(\text{sign}(y))$. Here $\text{sign}(y)$ is +1+1 for $y\geq 0$ and 1-1 otherwise, and the bin function maps $\{-1,+1\}$ back to $\{1,0\}$. The syndrome is then computed as $s=Hy_{b}^{\top}$. An error is detected if the syndrome is a non-zero vector (i.e., $s(y)\neq 0$). Following the pre-processing technique proposed by [bennatan2018deep], the input vector to the neural network is $[|y|,s(y)]$ with length n+(nk)n+(n-k) to avoid overfitting.
Diffusion Generative Models. Diffusion Models (DMs) [ho2020denoising, song2019generative, song2020score] are generative models that generate samples from a target data distribution, $p_{\text{data}}(x_{0})$ by reversing a predefined forward noising process [sohl2015deep]. In the forward diffusion process, a data sample $x_{0}$ is gradually perturbed with Gaussian noise over a continuous time interval $t\in[0,T]$. This forward process can be mathematically described as adding noise to obtain a noisy data point $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathit{\bm{I}})$ is standard Gaussian noise and $\alpha_{t}\in[0,1]$ monotonically decreases with time step $t$ to control the noise level. Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising] $\bm{\epsilon}_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\times[T]\mapsto\mathcal{X}$ is trained to predict the noise $\bm{\epsilon}_{t}$ at each time step $t$, also learn the score function of $p_{t}(\mathbf{x_{t}})$ [song2019generative, song2020score]:
$\min_{\theta}\mathbb{E}_{\mathbf{x}_{t},\bm{\epsilon}_{t},t}\left[\left\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{t}\right\|_{2}^{2}\right]=\min_{\theta}\mathbb{E}_{\mathbf{x}_{t},\bm{\epsilon}_{t},t}\left[\left\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)+\sqrt{1-\alpha_{t}}\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}_{\text{Score Function}}\right\|_{2}^{2}\right],$ (1)
During inference, samples can be generated by solving the reverse-time SDE starting from $t=T$ to $t=0$. Crucially, there exists a corresponding deterministic process, the probability flow ODE (PF-ODE), whose trajectories share the same marginal distributions ${p_{t}(x_{t})}_{t\in[0,T]}$ as the SDE [song2020score]. The formulation of PF-ODE can be described and simplified as following [karras2022elucidating, song2023consistency]:
$\mathrm{d}\mathbf{x}_{t}=-\dot{\sigma}(t)\sigma(t)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\mathrm{d}t=\frac{\mathbf{x}_{t}-\epsilon_{\theta}(\mathbf{x}_{t},t)}{t},$ (2)
Sample generation of diffusion models is performed by solving this PF-ODE backwards in time from $t=T$ to $t=0$, starting from a sample drawn from the prior Gaussian distribution, $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma(T)^{2}I)$. This requires a numerical ODE solver (e.g., Euler [song2019generative, song2020score] or Heun [karras2022elucidating]) to obtain a solution trajectory $\{\hat{\mathbf{x}}_{t}\}_{t\in[0,T]}$ that transforms noise into a data sample.
Denoising Diffusion Error Correction Codes. The application of diffusion models to error correction was pioneered by DDECC [choukroundenoising]. Its core insight is to model the transmission of a BPSK-modulated codeword $\mathbf{x}_{0}\in\{-1,1\}^{n}$ over an AWGN channel as the forward diffusion process. A received signal $y$ is treated as a noisy sample $\mathbf{x}_{t}$ at a specific timestep $t$, where the noise schedule is designed to match the channel’s characteristics. This forward process is described as:
$y\mathrel{\mathop{\ordinarycolon}}=\mathbf{x}_{t}=\mathbf{x}_{0}+\sqrt{\bar{\beta_{t}}}\epsilon,$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$, and the cumulative noise variance $\bar{\beta_{t}}=\sum_{i=1}^{t}\beta_{i}$ corresponds to the channel’s noise level $\sigma^{2}$. Decoding is then performed via an iterative reverse denoising process, starting with the received signal $y\mathrel{\mathop{\ordinarycolon}}=\mathbf{x}_{t}$ and applying the denoising update rule for multiple steps, with a trained denoising network $\epsilon_{\theta}(\cdot,\cdot)$ predicting the multiplicative noise:
$\mathbf{x}_{t-1}=\mathbf{x}_{t}-\frac{\sqrt{\bar{\beta}_{t}}\beta_{t}}{\bar{\beta}_{t}+\beta_{t}}(\mathbf{x}_{t}-\text{sign}(\mathbf{x}_{t})\epsilon_{\theta}(\mathbf{x}_{t},t)),$ (4)
A key innovation in DDECC is its conditioning mechanism. In the ECC domain, the number of parity check errors (syndrome sum), $e_{t}=\sum_{i}s(y)_{i}$, serves as a direct measure of the noisy level in a received signal $y$. Therefore, DDECC conditions its denoising network $\epsilon_{\theta}$ on the sum of syndrome error $e_{t}$ instead of a timestep $t$, making the diffusion models adapted to the structure of the error correction problem. The denoising network is trained to learn the hard prediction of the multiplicative noise with a Binary Cross-Entropy (BCE) loss:
$\mathcal{L}(\theta)=-\mathbb{E}_{e_{t},\mathbf{x}_{0},\epsilon}\log(\epsilon_{\theta}(\mathbf{x}_{0}+\sqrt{\bar{\beta}_{t}}\epsilon,e_{t}),\tilde{\epsilon}_{b}),$ (5)
where $\tilde{\epsilon}_{b}=\text{bin}(\mathbf{x}_{0}(\mathbf{x}_{0}+\sqrt{\bar{\beta}_{t}}\epsilon))$ denotes the target binary multiplicative noise.
4 Method: Consistency Flow Model for One-step decoding
4.1 Error Correction Consistency Flow Property
@@
Despite its impressive decoding performance, a drawback of the DDECC algorithm is the computational overhead during inference due to its iterative denoising mechanism. Inspired by the recent success of Consistency Models (CMs) [song2023consistency, song2023improved] in image generation, we propose the first approach for training consistency models for Error Correction Codes (ECC). Consistency Models (CMs) [song2023consistency] were introduced to overcome the inference computational costs by enabling fast, one-step generation. The core principle is that: any two points $(\mathbf{x}_{t},t)$ and $(\mathbf{x}_{r},r)$ on the same PF-ODE trajectory should map to the same origin point $\mathbf{x}_{0}$. CMs build upon Eq. 2 and learn a function $f_{\theta}(\mathbf{x}_{t},t)$ that directly estimates the trajectory from noisy data to clean data with a single step:
@@
$\mathcal{L}_{\text{Standard-CM}}(\theta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[w(t)d\left(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r)\right)],$ (7)
Here, $f_{\theta}$ is the consistency network being trained and $w(t)$ denotes the time schedule. This function is defined by two fundamental properties proposed by [song2023consistency, song2023improved]: 1) Boundary Condition: At timestep $t=0$, the function is the identity: $f_{\theta}(\mathbf{x}_{0},0)=\mathbf{x}_{0}$. 2) Self-Consistency: For any two points $(\mathbf{x}_{t},t)$ and $(\mathbf{x}_{r},r)$ on the same trajectory, the function yields the same output: $f_{\theta}(\mathbf{x}_{t},t)=f_{\theta}(\mathbf{x}_{r},r)$.
Unlike generation tasks, which aim to sample from a target data distribution $p_{\text{data}}$, the goal in ECC is to decode a single, ground-truth codeword $\mathbf{x}_{0}$ corresponding to a given received noisy signal $y$. The ideal target distribution for the decoder is therefore a point mass at the correct codeword, which we model conceptually as a Dirac delta function, $\delta(\mathbf{x}-\mathbf{x}_{0})$. Based on this, we propose that a consistency-based decoder should satisfy the following Error Correction Consistency Flow condition:
Error Correction Self-consistency. Given a trajectory $\{y\mathrel{\mathop{\ordinarycolon}}=\mathbf{x}_{t}\}_{t\in[0,T]}$, we learn the consistency function as $f\mathrel{\mathop{\ordinarycolon}}(\mathbf{x}_{t},t;\theta)\mapsto\mathbf{x}_{0}$, holding the Error Correction Consistency Flow property: for a given ground-truth codeword $\mathbf{x}_{0}$, all points $(\mathbf{x}_{t},t)$ on any trajectory originating from $\mathbf{x}_{0}$ map directly back to it, i.e., $f_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{0},\quad\forall t\in[0,T]$. This implies that for any two noisy signals $\mathbf{x}_{t}$ and $\mathbf{x}_{r}$ derived from the same $\mathbf{x}_{0}$, their consistency function outputs must be identical and correct: $f_{\theta}(\mathbf{x}_{t},t)=f_{\theta}(\mathbf{x}_{r},r)=\mathbf{x}_{0},\forall t,r\in[0,T]$. We learn the consistency model such that the conditional distribution is modeled as $p_{\theta}(\mathbf{x}|\mathbf{x}_{t})=\delta(\mathbf{x}-f_{\theta}(\mathbf{x}_{t},t))$, where $f_{\theta}(\mathbf{x}_{t},t)$ directly estimates $\mathbf{x}_{0}$.
To train a consistency model $f_{\theta}$ that achieves this property, a naive adaptation of the vanilla consistency loss would be to minimize the distance between two noisy signals from the same trajectory similar to Eq. 7. However, this standard consistency objective is indirect; it only enforces relative consistency between outputs. In decoding tasks, the ground truth $\mathbf{x}_{0}$ is known during training, and consistency models are optimized to decode the clean codeword. We leverage this property by proposing the Error Correction Consistency Flow Loss (EC-CM), which minimizes the distance of each estimation to the ground truth $\mathbf{x}_{0}$:

### v2

# Syndrome-Flow Consistency Model Achieves One-step Denoising Error Correction Codes

@@ Abstract
Error Correction Codes (ECC) are fundamental to reliable digital communication, yet designing neural decoders that are both accurate and computationally efficient remains challenging. Recent denoising diffusion decoders achieve state-of-the-art performance, but their iterative sampling limits practicality in low-latency settings. To bridge this gap, consistency models (CMs) offer a potential path to high-fidelity one-step decoding. However, applying CMs to ECC presents a significant challenge: the discrete nature of error correction means the decoding trajectory is highly non-smooth, making it incompatible with a simple continuous timestep parameterization. To address this, we re-parameterize the reverse Probability Flow Ordinary Differential Equation (PF-ODE) by soft-syndrome condition, providing a smooth trajectory of signal corruption. Building on this, we propose the Error Correction Syndrome-Flow Consistency Model (ECCFM), a model-agnostic framework designed specifically for ECC task, ensuring the model learns a smooth trajectory from any noisy signal directly to the original codeword in a single step. Across multiple benchmarks, ECCFM attains lower bit-error-rate (BER) and frame-error-rate (FER) than transformer-based decoders, while delivering inference speeds 30x to 100x faster than iterative denoising diffusion decoders.
@@ 1 Introduction
Error Correction Codes (ECC) play a central role in modern digital communications and have been essential in a wide range of applications, including wireless communication and data storage. The core task of an ECC decoder is to recover a message from a received signal corrupted by noise during transmission. Recently, inspired by the great success of deep neural networks, neural network-based decoders were introduced, which are capable of improving the performance scores of conventional, problem-specific algorithms such as Belief Propagation (BP) [33] and Min-Sum (MS) [11].
@@
Figure 1: Illustration of our proposed Error Correction Syndrome-Flow Consistency Model (ECCFM). ECCFM learns to map received signals from the trajectories to a single, consistent codeword prediction $\mathbf{x}_{0}$, represented by a $\delta$ function.
The existing neural decoders can be categorized into two groups. Early model-based approaches [25, 28, 43] achieved successful results by integrating neural networks into conventional decoding algorithms. However, their reliance on problem-specific structures can limit the applicability as a general-purpose decoder. To address this limitation, model-free decoders enable the extension of a general neural network architecture without prior knowledge of decoding algorithms. While the early proposals applying fully connected neural net architectures [13, 3, 17] could lead to overfitting, the preprocessing techniques that utilize magnitude and syndrome vectors [1] have been effective in reducing the error. These improvements have led to efficient transformer-based decoders [5], which leverage the self-attention mechanism to improve the numerical performance on short block codes. The transformer-based decoders have been improved in several recent works [7, 8]. Notably, the recent work by [31] integrates cross-attention between magnitude and syndrome, resulting in the state-of-the-art performance in the decoding task.
@@
In parallel, the remarkable success of diffusion generative models [15, 40] has inspired a new direction for neural decoder. Diffusion models train a noise estimator to gradually denoise noisy input data and reverse a forward noise manipulation process, generating high-quality codewords by an iterative denoising process. DDECC [4] first extends the denoising diffusion framework to ECC, naturally modeling the AWGN channel as a forward diffusion process. A time-dependent transformer learns to denoise the received signal iteratively, recovering the original codeword. This framework provides complementary gains over transformer-based methods, particularly for long codes and low SNR, due to its iterative refinement denoising process.
However, the iterative denoising process inherently suffers from high computational complexity and latency, creating a significant bottleneck for real-time communication systems. While the iterative refinement is the key to high accuracy, it stands in direct conflict with the requirement for fast decoding in practical deployments. This challenge prompts the following question: Can we design a neural decoder that maintains the performance of diffusion-based decoders but with significantly reduced latency?
To solve this tension between accuracy and latency, the consistency model (CM) framework proposed by [37] for general applications provides a potential path. Designed to directly learn the reverse denoising trajectory into a single, efficient step, CMs could potentially offer a direct path to a high-fidelity, one-step decoder. However, our numerical results indicate that a direct application of existing CMs to ECC decoding would not yield the optimal performance, as detailed in Appendix 7.5.7. This outcome appears to arise from a conceptual incompatibility between the continuous nature of CMs and the discrete nature of the ECC task. As we discuss, the core assumption of a standard CM is the existence of a smooth trajectory along a Probability Flow ODE (PF-ODE). In ECC, the level of corruption of a valid codeword is measured by sum of syndrome errors [6]. The decoding path would not be a smooth flow but a series of jumps across a discrete space of error patterns, leading to suboptimal results.
Our main contribution is to resolve this incompatibility by re-parameterizing the decoding trajectory itself. Following the recent work by [20], we introduce a soft-syndrome as a time condition instead of the standard syndrome, which provides a continuous and differentiable measure of the level of corruption of signals. By using the soft-syndrome condition to parameterize the reverse PF-ODE process, we build a smooth trajectory that is native to the ECC problem, mapping the noisy signal towards the manifold of clean codewords. Building on this, we propose the Error Correction Syndrome-Flow Consistency Model (ECCFM), a model-agnostic consistency framework for training single-step decoders. The ECCFM is trained with a well-designed consistency objective that maps any points on the same soft-syndrome-defined trajectory to the original clean codeword, as illustrated in Figure 1. This forces the model to learn the decoding trajectory by a single consistency function.
@@
We numerically evaluate the effectiveness of ECCFM through experiments on a diverse set of standard codes, including BCH, LDPC, Polar, MacKay, and CCSDS. The results show that ECCFM can improve bit-error-rate (BER) and frame-error-rate (FER) compared to leading transformer-based methods, exhibiting strong gains on longer codes with code lengths above $n=200$. Notably, it achieves an inference speedup of 30x to 100x over the existing iterative denoising diffusion methods in a single step, while maintaining comparable decoding performance. In our numerical evaluation, ECCFM offers a model-free training paradigm for building ECC decoders that achieves near state-of-the-art performance with low latency, and requires only a single-step inference suited for real-world communication systems.
@@ 2 Related Works
Neural Network-based ECC Decoders. Neural decoders are broadly categorized into model-based and model-free approaches. Model-based methods unfold and enhance conventional algorithms like Belief Propagation (BP) [33] and Min-Sum (MS) [11] with learnable message-passing components. This paradigm, explored across various architectures [9, 18, 19, 25, 28, 29, 26, 43], consistently surpasses its traditional counterparts in performance [27], but can be constrained by the underlying algorithm’s structure and difficulties in capturing long-range dependencies.
@@
In contrast, model-free decoders treat decoding as an end-to-end learning problem. Early fully-connected architectures [13, 3, 17] suffered from overfitting, a challenge later mitigated by pre-processing techniques like magnitude-syndrome decomposition [1]. The advent of Transformers [42] led to ECCT [5], which framed decoding as an auto-regressive sequence-to-sequence task. Subsequent works built upon this foundation to improve generalization (FECCT [7]), enable joint encoder-decoder training (DC-ECCT [8]), and achieve superior performance among auto-regressive methods with cross-attention (CrossMPT [31]). More recently, DDECC [4] introduced a diffusion-based approach, framing decoding as a denoising process and outperforming auto-regressive decoders.
Diffusion Generative Models. Diffusion generative models [35, 15, 40, 16] have achieved state-of-the-art synthesis quality across diverse domains such as images [10, 34, 32], video [14, 2], text [23, 30], and graphs [41, 22, 21]. Their primary limitation, however, is the significant inference latency caused by the iterative denoising process, which has spurred the development of numerous accelerated sampling methods [36, 24].
The diffusion framework is particularly well-suited for ECC, as the AWGN channel provides a forward process [4]. While diffusion-based decoders have set a new state-of-the-art in performance [4, 31], they inherit the same critical issue of high computational cost. To resolve this accuracy-latency trade-off, we turn to Consistency Models (CMs) [37, 38, 12]. CMs are a recent advance in diffusion model acceleration, designed to learn a direct mapping from any point on a denoising trajectory to its origin. This allows for high-fidelity, single-step generation, offering a potential path toward efficient and powerful ECC decoders.
@@ 3 Preliminaries
The objective of a decoder is to estimate the original codeword $\hat{x}$ from the noisy signal $y$. An essential tool for error detection is the syndrome, calculated from a hard-demodulated formulation of the received signal, $y_{b}$, performed as $y_{b}=\text{bin}(\text{sign}(y))$. Here $\text{sign}(y)$ is +1+1 for $y\geq 0$ and 1-1 otherwise, and the bin function maps $\{-1,+1\}$ back to $\{1,0\}$. The syndrome is then computed as $s=Hy_{b}^{\top}$. An error is detected if the syndrome is a non-zero vector (i.e., $s(y)\neq 0$). Following the pre-processing technique proposed by [1], the input vector to the neural network is $[|y|,s(y)]$ with length n+(nk)n+(n-k) to avoid overfitting.
@@
Consistency Models. CMs [37] built upon diffusion generative models and were introduced to overcome the inference computational costs by enabling fast, one-step generation. The core principle is that: any two points $(\mathbf{x}_{t},t)$ and $(\mathbf{x}_{r},r)$ on the same PF-ODE trajectory should map to the same origin point $\mathbf{x}_{0}$. CMs learn a function $f_{\theta}(\mathbf{x}_{t},t)$ that directly estimates the trajectory from noisy data to clean data with a single step:
@@
$\mathcal{L}_{\text{Standard-CM}}(\theta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[w(t)d\left(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r)\right)].$ (2)
Here, $f_{\theta}$ is the consistency network being trained and $w(t)$ denotes the time schedule. We provide a detailed review of diffusion generative models and consistency models in the Appendix, Section 7.1.
4 Methodology
4.1 Syndrome-Flow Consistency Property
((a))
While diffusion-based decoders like DDECC [6] achieve high performance, their iterative nature imposes significant computational overhead at inference. To enable fast, single-step decoding, we move beyond iterative refinement and propose a new paradigm based on the Syndrome-Flow Consistency, which defines the behavior of a one-step decoder. The consistency decoder should learn a function, $f_{\theta}$, that deterministically maps any noisy signals, $\mathbf{x}_{t}$, directly to its clean origin, $\mathbf{x}_{0}$, regardless of the noise level $t$. This concept is inspired by the self-consistency principles of Consistency Models (CMs) [37, 38], but is specifically tailored to the deterministic nature of the error correction task. In decoding, the goal is not to sample from a distribution, but to recover a single, correct codeword defined by a Dirac delta function $\delta(\mathbf{x}-\mathbf{x}_{0})$. Formally, we define the Syndrome-Flow Consistency property as follows:
Syndrome-Flow Consistency Property. Given a trajectory $\{y\mathrel{\mathop{\ordinarycolon}}=\mathbf{x}_{t}\}_{t\in[0,T]}$, we learn the consistency function as $f\mathrel{\mathop{\ordinarycolon}}(\mathbf{x}_{t},t;\theta)\mapsto\mathbf{x}_{0}$, holding the Syndrome-Flow Consistency property: for a given ground-truth codeword $\mathbf{x}_{0}$, all points $(\mathbf{x}_{t},t)$ on any trajectory originating from $\mathbf{x}_{0}$ map directly back to it, i.e., $f_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{0},\quad\forall t\in[0,T]$. This implies that for any two noisy signals $\mathbf{x}_{t}$ and $\mathbf{x}_{r}$ derived from the same $\mathbf{x}_{0}$, their consistency function outputs must be identical and correct: $f_{\theta}(\mathbf{x}_{t},t)=f_{\theta}(\mathbf{x}_{r},r)=\mathbf{x}_{0},\forall t,r\in[0,T]$. We learn the consistency model such that the conditional distribution is modeled as $p_{\theta}(\mathbf{x}|\mathbf{x}_{t})=\delta(\mathbf{x}-f_{\theta}(\mathbf{x}_{t},t))$, where $f_{\theta}(\mathbf{x}_{t},t)$ directly estimates $\mathbf{x}_{0}$.
A naive application of the standard consistency training objective (Eq. 2) is insufficient, as it only enforces relative consistency by minimizing $d(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r))$. This is an indirect and inefficient objection for ECC task, where the ground truth $\mathbf{x}_{0}$ is known during training. We therefore introduce the Error Correction Consistency Loss (EC-CM), which directly enforces our proposed property by minimizing the distance between each predicted outcome and the true codeword $\mathbf{x}_{0}$: